OpenStudy (nottim):

Question 8 of Integration Calculus

OpenStudy (nottim):

Anyone one would do.

OpenStudy (science0229):

Hello

OpenStudy (science0229):

Let's start with part a

OpenStudy (science0229):

The problem made a substitution of u=x^8+12, right?

OpenStudy (science0229):

@NotTim

OpenStudy (nottim):

eh, i thinkso

OpenStudy (science0229):

Now, can you find du?

OpenStudy (nottim):

8x^7

OpenStudy (science0229):

Right.

OpenStudy (nottim):

7/8(x^8+12)

OpenStudy (science0229):

Wait. Let me rephrase the question. Sorry.

OpenStudy (science0229):

Forget what I just asked.

OpenStudy (science0229):

You got to the point where du=8x^7 What number do you have to multiply by 8x^7 in order to get to 7x^7?

OpenStudy (nottim):

7/8?

OpenStudy (science0229):

Right!

OpenStudy (science0229):

Now, the original problem was \[\int\limits_{}^{}7x^7 \sqrt{x^8+12}dx=\int\limits_{}^{}A \sqrt{u}du\]

OpenStudy (science0229):

We agreed that du=8x^7

OpenStudy (nottim):

ja

OpenStudy (nottim):

ya

OpenStudy (nottim):

i gotta go; i'll be back real soon.

OpenStudy (science0229):

I have to go, too...

OpenStudy (nottim):

oh darn

OpenStudy (nottim):

???